Which answer correctly shows the first step that can be used to correctly solve the system of equations?
5x + 2y = 4
3x + 4y + 2z = 6
7x + 3y + 4z = 29

A.
3x + 4y + 2z = 6 → 3x + 4y + 4z = 6
7x + 3y + 4z = 29 → (­-)7x + 3y + 4z = 29
-4x ­- y =- ­23
B.
3x + 4y + 2z = 6 → 6x + 8y + 4z = 12
7x + 3y + 4z = 29 → (-­)7x + 3y + 4z = 29
-x + 5y = ­-17
C.
3x + 4y + 2z = 6 → 6x + 8y + 4z = 6
7x + 3y + 4z = 29 → (-­)7x + 3y + 4z = 2
-x + 5y =- ­23
D.
3x + 4y + 2z = 6 → 6x + 8y + 4z = 12
7x + 3y + 4z = 29 → (-­)7x + 3y + 4z = 29
-x + 8y = 41
Thanks in advance, this has been a difficult problem for me.

Answer by Edwin McCravy(20054)

5x + 2y = 4 
3x + 4y + 2z = 6 
7x + 3y + 4z = 29


A.	
3x + 4y + 2z = 6  →     3x + 4y + 4z = 6 
Oh,oh! They forgot to multiply the 6 by 2>
7x + 3y + 4z = 29 → (­-)7x + 3y + 4z = 29

-4x ­-  y  =- ­23

B.	
3x + 4y + 2z = 6 →     6x + 8y + 4z = 12 
 7x + 3y + 4z = 29 → (-­)7x + 3y + 4z = 29

-x + 5y  = ­-17

C.	
3x + 4y + 2z = 6 →     6x + 8y + 4z =  6
Oh,oh! They forgot to multiply the 6 by 2>
7x + 3y + 4z = 29 → (-­)7x + 3y + 4z = 2

-x + 5y =- ­23

D.	
3x + 4y + 2z = 6 →     6x + 8y + 4z = 12 
7x + 3y + 4z = 29 → (-­)7x + 3y + 4z = 29

-x + 8y = 41
Oh,oh! They didn't subtract 3y from 8y,
and also they added 29 to 12 when they should 
have subtracted.>

Edwin